A force P is applied at B, in a regular hexagon, as shown. Determine

 (a) the moment of the force P about E, 

 (b) the magnitude and sense of the horizontal force applied at A that creates the same moment about E, 

 (c) the (i) magnitude and (ii) direction (from horizontal) of the smallest force applied at A that creates the same moment about E. (Hint: Internal angle of regular hexagon is 120°) 

 

 If P=500 N, a=40°, a=300 mm
Answer #1

600 - (a) from diagram, X+60+ 8 = 180 B = 120-2 Nos, perb. seberation between fosce P & point E A LE = X = BE Sim B → X=2a Sihorizontally. (b). let force (P) on A acts LCDE = 1200 LL DE = 60° - DE Perpendicularo seberation betworen / force P & point(C) tet force () on A acts at angle o from horizontal. from diagram, AE: 3 a A A 28-12 T 20 BC f LLAE = (@+30) Nag In A ALE,2954.4 2954.4 *** = 568.59 N-m P Sin (0+30) = - 353 5.196 nax. or min, p Sin CO +30) - 568.59 a pe 568.59 Corec (0+30) - 0 fo
Answer #2

Not sure if the forces around A are what this question is asking for, but are they correct for what they are?  How does the moment around E look?  Thanks for any advice.  :)DDA2B5FC-F5D0-47A5-BFCB-321171C5BEE9.jpeg

source: Me
answered by: Chris